zaky21:

A solenoid has 450 loops, each of radius 0.0254m. the field increases from0 t to 3.00 T in 1.55 s. what is the EMF generated in the coil? (Hint: what is the area of a circle?) (Unit=volts)

TheSmartOne:

Hi there! So do you know Faraday's law? That's what it looks like we'll need to use How do you calculate the area of a circle?

TheSmartOne:

@zaky21 are you still there?

zaky21:

yes

zaky21:

yes

TheSmartOne:

So what's Faraday's law? Also what's the formula for finding the area of a circle?

zaky21:

A=πr2

TheSmartOne:

So what is the area if the radius is 0.0254 m

zaky21:

2.03×10-3

TheSmartOne:

Good! Now the formula we need to be using is EMF = -n * \(\frac{\Delta (BA)}{\Delta t}\)

TheSmartOne:

We just found A, which is the area We are given what B is. It goes from 0 to 3.00 T So what is the change in the magnetic field? It would be the final - initial So what would that be?

zaky21:

i dont get it

TheSmartOne:

So the formula has \(\Delta\) That means the change in Our magnetic field went from 0 to 3 0 is our initial, 3 is our final magnetic field So the change in the magnetic field is final minus initial What would that be? Just told you which number is initial and final as well

zaky21:

so wat do i subtract

TheSmartOne:

0 is initial 3 is final Do final - initial

TheSmartOne:

Final minus initial

zaky21:

so 0-3

zaky21:

so 0-3

zaky21:

so 0-3

TheSmartOne:

Is 0 the initial value or the final value?

zaky21:

initial?

TheSmartOne:

Yes so are we doing final - initial Or initial - final?

zaky21:

initial

TheSmartOne:

That doesn't answer what I asked Should it be 3-0=? Or should we be doing 0-3=?

zaky21:

0-3

TheSmartOne:

bro you're so hopeless FINAL MINUS INITIAL 3 IS FINAL 0 IS INITIAL SO DO FINAL MINUS INTIAL WHAT DO YOU GET

zaky21:

its 3

zaky21:

@TheSmartOne

zaky21:

its 3 bro

TheSmartOne:

Yes now plug it all in

TheSmartOne:

EMF = -n * \(\frac{\Delta (BA)}{\Delta t}\) n = 450 B = 3 A = 2.03*10^-3 t = 1.55

zaky21:

−1.76806451 thats wat i got

TheSmartOne:

Yes

zaky21:

i have another one

TheSmartOne:

New post then

zaky21:

i did